a)     A simple closed economy with an mpc equal to 0.5. Investment spending has suddenly fallen, reducing aggregate demand and output to a level that is 100 million below Y*.

                   i.           If the government decide to try to get the economy back to full employment using only an increase in government spending, by how much would G need to be increased?

                 ii.           If the government, instead, decided to try to get the economy to full employment using only a lump-sum tax cut how big of a tax cut would be needed?

      


1
Expert's answer
2021-07-06T09:01:32-0400

(i)

"Multiplier=\frac{1}{1-0.5}"

"∆Y=mult\times∆G"

"100 million=2\times∆G"

"\u2206G=50million"

The government will have to increase spending by 50 million.

(ii)

Tax multiplier, -mpc (mult)

"∆Y=-mpc(mult)\times∆T"

"∆Y=-0.5\times2×∆T"

"∆T=\frac{100million}{-1}"

"\u2206T=-100 million"

The government will have to give a lump-sum tax cut of 100 million.